solve the following equations;
3x+2y+7=0
6x+4y-8=0,
a. algebraically
b. graphically

Answer by mananth(16949)

3 x + 2 y = -7 .............1
6 x + 4 y = 8 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-6 x -4 y = 14
6 x + 4 y = 8
Add the two equations
0 x = 22
0=22
NO solution
The lines are parallel


equation1
when x= 0 y= -3.5 ( 0 , -3.5 )
when y = 0 x= -2.33 ( -2.33 , 0 )
Equation 2
when x= 0 y= 2 ( 0 , 2 )
when y = 0 x= 1.33 ( 1.33 , 0 )

m.ananth@hotmail.ca